OpenStudy (anonymous):

I am taking Algebra 2 and am completely hopeless..can anyone explain the transformations of quadratic functions????

OpenStudy (anonymous):

y= a (x-h)^2 + k h shift to the right k shift to the up/down a determine the width

OpenStudy (anonymous):

let's to an example start with something simple y= x^2 y= (x-2)^2+3 , shifts right by 2, up by 3 y= (x+3)^2+5 , shift left by 3, up by 5

OpenStudy (anonymous):

written like that is vertex form,so..okay,is the "h" always going to be left or right depending on if its positive or negative?

OpenStudy (anonymous):

well if h is positive , it shifts right

OpenStudy (anonymous):

okay,so what about compressions and stretches?

OpenStudy (anonymous):

a does that

OpenStudy (anonymous):

bigger than a , narrower the parabola